OpenStudy (anonymous):

when isabella pushes a box of mass 5.25 kilograms with a force of 15.7 newtons, it accelerates by a rate of 2.50 meters/second. what is the force due to friction in this scenario?

OpenStudy (anonymous):

let force of friction is f so 15.7-f=5.25*2.50

OpenStudy (anonymous):

also unit of acceleration is m/s^2

OpenStudy (theeric):

Hello, @dirtslinger28 , do you have any more questions on this one? You see, \(F_{net}=ma\) But you don't know \(F_{net}\). And \(F_{net}=F_{push}+F_{friction}\) So what you're looking at now is that you know everything except for \(F_{net}\) and \(F_{friction}\). You can solve this problem by substituting for \(F_{net}\). Then \(F_{friction}\) is the only unknown variable. \(F_{net}=F_{push}+F_{friction}=ma\) We look at the right two expressions, now. \(F_{push}+F_{friction}=ma\) That's where gyanu left you. Now you can solve for \(F_{friction}\), plug in your numbers, and calculate.

OpenStudy (anonymous):

F = ma Force is in Newtons and you have a known 15.7N pushing the box 15.7N = 5.25 kg * acceleration a = 15.7 / 5.25 = 2.99 m/s^2 So the acceleration with no friction would be 2.99 m/s^2 Since your box goes 2.5 m/s^2, that means the friction force is causing the box to slow by 2.99 m/s^2 - 2.5 = 0.49 m/s^2 Friction is a force and not an acceleration so you have to convert that to Newtons for the friction force: F = ma F = 5.25kg * 0.49 m/s^2 F = 2.6 Newtons , is lost to friction

OpenStudy (theeric):

biire2u makes use of the interesting concept of acceleration being additive :)